OpenStudy (annjewell1998):

If y varies directly with x and y = -7 when x = -1, find x when y = -84. a. 84 b. 7 c. -12

OpenStudy (princeharryyy):

let y=k*x where k is some constant the, -7 = k*(-1) substituting in above equation. => k = 7 there fore the equation y=k*x becomes y=7x put y =-84 => -84 = 7*x =>x=-12

OpenStudy (annjewell1998):

Oh okay, thank you!!

OpenStudy (princeharryyy):

:)